OpenStudy (anonymous):

what's the slant height of this pyramid whats the lateral area total area volume

OpenStudy (anonymous):

@jim_thompson5910 can u help?

jimthompson5910 (jim_thompson5910):

the base of the red dashed right triangle is 5 (since this is half of 10). The height of the red dashed triangle is 12. Let's use these facts to find the hypotenuse of the red dashed triangle (which is the slant height). a^2 + b^2 = c^2 5^2 + 12^2 = c^2 25 + 144 = c^2 169 = c^2 c^2 = 169 c = sqrt(169) c = 13 So the slant height is 13 units

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

okay:) yes!!

jimthompson5910 (jim_thompson5910):

alright now onto the surface area

jimthompson5910 (jim_thompson5910):

the area of one lateral face is A = (bh)/2 A = (10*13)/2 A = 130/2 A = 65 So the area of one lateral face is 65 square units So the area of all four lateral faces is 65*4 = 260 square units

jimthompson5910 (jim_thompson5910):

So the total lateral area is 260 square units

OpenStudy (anonymous):

got it!

jimthompson5910 (jim_thompson5910):

The area of the base is A = s^2 = 10^2 = 100 So add this to the lateral area: 260+100 = 360 Therefore the total surface area is 360 square units

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

Finally, let's find the volume

jimthompson5910 (jim_thompson5910):

V = (Area of base*Height)/3 V = (100*12)/3 V = 1200/3 V = 400 So the volume is 400 cubic units

jimthompson5910 (jim_thompson5910):

Hopefully that all makes sense

OpenStudy (anonymous):

thank you soo much!!:D

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

glad I could help out again